There is a two-digit number whose digits are the same, and has got the following property: When squared, it produces a four-digit number, whose first two digits are the same and equal to the original’s minus one, and whose last two digits are the same and equal to the half of the original’s. Find that number.
Answer by josgarithmetic(39617)

------------------------------------------------------------
two-digit number whose digits are the same, and has got the following property: When squared, it produces a four-digit number, whose first two digits are the same and equal to the original’s minus one, and whose last two digits are the same and equal to the half of the original’s.
------------------------------------------------------------

22*22 is a three-digit number.
33*33 is a four-digit number.
The two-digit number must be 33 or greater.

33^2=1089
44^2=1936
55^2=3025
66^2=4356
77^2=5929
88^2=7744
99^2=9801

Description seems closest to 88 as the original number.